OpenStudy (anonymous):

There are two fruit trees located at (3,0) and (–3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane. Can someone please help? I really want to know how to do this but it is hard to understand. I will give a medal!!!

OpenStudy (anonymous):

you can graph the equations using desmos

OpenStudy (anonymous):

thanks for replying. but honestly, Im doing this online and I don't understand anything. I need someone to help me step by step

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

yes im here. give me a second

OpenStudy (anonymous):

np just making sure

OpenStudy (anonymous):

okay. what else do I do?

OpenStudy (anonymous):

does it look like this?

OpenStudy (anonymous):

it can't have a zero.... sorry i completely misread your problem, lets start over:

OpenStudy (anonymous):

the points are focal points, not vertexes, so it won't work

OpenStudy (anonymous):

okay. sure :)

OpenStudy (anonymous):

you can use :a^2−b^2=c^2 3,4,5 work to get: a=5,b=4 c=3 and use x^2/5^2+y^2/4^2=1 this graph will make your foci (−3,0) and (3,0)

OpenStudy (anonymous):

i just attached the graph

OpenStudy (anonymous):

for the hyperbola x^2/a^2−y^2/b^2=1 you can use this formula: a^2+b^2=3^3 make a^2=8,b^2=1 to get 9, then plug them in to get: x^2/8−y^2=1

OpenStudy (anonymous):

x^2/8−y^2=1 this is your hyperbola

OpenStudy (anonymous):

if you graph the hyperbola, ellipse, and two points, on one graph you will have your answer

OpenStudy (anonymous):

ooh okay. thanks, ill do it then show you what I have

OpenStudy (anonymous):

is what it should look like

OpenStudy (anonymous):

sorry it took a bit long and i messed up at first

OpenStudy (anonymous):

thats okay, thanks